Project

General

Profile

Actions

Agenda

1 Sign off from equipment owners
  • John Cobb (FC)
  • Alan Bross (SSU/SSD)
2 List of process variables to monitor on the Focus Coil
  • where are they in the GUI?
  • what are the low and high limits?
  • what is the monitoring frequency?
3 List of process variables to monitor on SSU and SSD
  • where are they in the GUI?
  • what are the low and high limits?
  • what is the monitoring frequency?

4 Who is on-call at a given time, for each magnet? What are their contact details, if they are not on the expert list?

5 Propose: the magnets to be ramped under DC control and handed back to DC control before powering down. The hall should be left in a "controlled access/controlled entry state"; mechanical beamstop removed and emergency stops opened, such that we can begin data taking without entering the hall.

I propose a discussion at 15:00 in the MICE meeting room to decide if these prerequisites have been met. We can use

http://mice.iit.edu/phonebridge.html

to decide if these prerequisites have been met.

Notes

Present:
Mike Courthold
Chris Rogers
Ken Long
Paolo Franchini
John Cobb
Josef Boehm
Victoria Blackmore
Pierrick Hanlet
Alan Bross

We started with discussion of water issues
  • Colin wanted to change pipes am 21/09/16. It did not appear to be a pressing need to change the piping. However the FC appears to have too high water temperature on the compressors. The delta T in the heat exchanger is too low. So the source for the water is the same - the chiller - regardless of the piping. John Govans is trying to get access to chiller to push the temperature down lower. Also could look at distribution of flow through the FC and increase flow through FC to improve the heat exchanged through FC. This is under investigation now (hall access ongoing).
  • If the water flow issue is not addressed, we should power the FC cautiously.
  • Note a small effect on He temperature (mK)
Pierrick
  • Concern that compressors are too close to the limit; Josef says no it is okay
Update on spectrometer
  • Colin put air to it; we have good flow, but not checked that leads were dry.
  • Spectrometer solenoids are stable.
  • Analogue ground fault detector needs to be logged to check for ground faults, as per Alan Bross email.
  • Operationally, we run with a risk of ground fault occurring with the trip removed, could endanger the magnet if there is a quench but no historical evidence of such.
Item 1: Sign off from equipment owners
  • John Cobb: no sign off
    • when/if water issues are resolved
  • Alan Bross: no sign off
    • when the QPS ground fault detector is evaluated as good
    • when the 8 PSI valves are reinstated (discussed later)
Item 2 and 3: Monitoring
  • As per table 1 of SSU document and table 1 of FC document
  • Note ignore the red "not suitable for combined monitoring" text in the document headers
  • Can the alarm handler be trusted? Pierrick: I believe it is reliable except for FC. i.e. NO.
  • Experts need to be happy with the alarm limits. Shifters will call the experts if the alarm handler goes off.
  • Write it down as well on 30 minute monitoring in the "MICE Control room log book"

We require strip tool to be open - Pierrick should provide

PSI Valves (back to Item 1)

  • Bross asked whether "8 PSI pressure relief valves" have been reinstated.
  • 8 PSI valves have not been reinstated; without that we cannot run. They are sitting in bits in the MLCR. They will need to be recalibrated.
    • Could work around by logging pressure; comment that it is too risky. It is a few hours work. We will investigate with Mark Tucker (Victoria Blackmore and Josef discussed with Mark Tucker in break out phone call).
    • Ken Long - who is responsible person? Josef, but Josef needs to then be included in the information dissemination and spectrometer needs to have change control. Note web whiteboard has lapsed. Josef is also the on-site person for the FC.
    • Ambitious to get two valves ready by today; 8 PSI "cracking pressure"
    • Following conversation with Mark Tucker, offline, should be installed in 1 hour.
On-call rota
  • Mike notes that Josef/Vicky/Mark Tucker are "competent experts". Mike notes other potentials are on the on-call list who could be promoted.
  • Decision: will call DC if there is an incident; DC is expected to disseminate to Josef and Pierrick
  • Victoria is DC (holds DC on-call phone).

Item 5

All agreed to the proposed operational handling.

Updated by Rogers, Chris about 7 years ago ยท 2 revisions